OpenStudy (anonymous):

A film set designer is using white and colored tiles in a pattern to create paths of different lengths. If x is the length of the path in feet, the number of colored tiles needed to make it is calculated using the rule (6x^2y^3)(4xy^3) where y represents the number of white tiles. Which simplified expression represents the number of colored tiles used for a path of length x feet?

OpenStudy (anonymous):

24x^2y^6 24x^3y^6 24x^2y^9 24x^3y^9

OpenStudy (anonymous):

Multiply them..

OpenStudy (anonymous):

my answer was b double checking

OpenStudy (anonymous):

\[\large (6x^2y^3)(4xy^3) = (6 \times 4)(x^2 \times x)(y^3 \times y^3) = ??\]

OpenStudy (anonymous):

Then you are right..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome..